# SIPDO: Closed-Loop Prompt Optimization via Synthetic Data Feedback

### v2

@@ Abstract
SIPDO: Closed-Loop Prompt Optimization via Synthetic Data Feedback
Yaoning Yu1 Ye Yu1 Kai Wei2 Haojing Luo3 Haohan Wang1 1University of Illinois at Urbana–Champaign 2University of South Florida 3 iDreamer.ai
Preprint. Under review.
@@ 1 Introduction
Large language models (LLMs) have demonstrated strong performance across a wide range of natural language tasks, including classification, question answering, and reasoning. However, their output quality is highly sensitive to prompt design—small changes in phrasing, structure, or formatting can lead to significant variations in performance He et al. (2024); Spiess et al. (2025). This sensitivity has made prompt optimization a core challenge in adapting LLMs to downstream applications.
@@
Prior work in prompt optimization has explored manual tuning, discrete search, and gradient-based methods to improve model responses Wang et al. (2023); Shin et al. (2020); Cui et al. (2024); Kwon et al. (2024); Zhang et al. (2024). While effective in some settings, these methods typically assume a fixed input distribution and treat prompt optimization as a one-time procedure. As a result, they may produce prompts that perform well on average but offer limited support for iterative improvement or adaptation to evolving failure modes.
In contrast, data augmentation is a long-standing technique in supervised learning for improving model robustness and generalization by exposing models to diverse training conditions Mikołajczyk and Grochowski (2018). In the context of prompt learning, the ability of LLMs to generate high-quality synthetic data opens the door to new forms of self-improvement. However, most existing prompt optimization pipelines do not leverage synthetic data in an adaptive or feedback-driven manner Singh et al. (2023); Gilardi et al. (2023); Tang et al. (2023); Gao et al. (2023).
@@
We propose SIPDO (Self-Improving Prompts through Data-Augmented Optimization), a closed-loop prompt optimization framework that integrates synthetic data generation directly into the learning process. SIPDO couples two components: a synthetic data generator that produces inputs designed to challenge the current prompt, and a prompt optimizer that uses these examples to refine the prompt iteratively. This feedback loop enables prompts to improve over time by addressing their own failures, without requiring access to new tasks or external supervision.
@@
  • We empirically demonstrate that augmenting prompt optimization with synthetic data improves performance across reasoning and QA benchmarks, surpassing existing prompt tuning methods.
@@ 2 Related Work
Automatic Prompt Engineering: Automatically discovering optimal prompts has become a key challenge in the era of large language models (LLMs). Automatic Prompt Engineering (APE) employs optimization-based, generative, and template-driven approaches. Optimization techniques include gradient-based search Shin et al. (2020), reinforcement learning Ouyang et al. (2022); Kwon et al. (2024), and evolutionary algorithms Cui et al. (2024). Generative methods use models like GPT and Gemini to generate candidate prompts, with StablePrompt Kwon et al. (2024) optimizing prompts via reinforcement learning. Additionally, PromptAgent Wang et al. (2023) breaks down prompt creation into sub-goals, while template-driven approaches, like fill-in-the-blank formats, ensure clarity Chen et al. (2024).
Recent work has expanded on automatic prompt optimization techniques. AutoPDL Spiess et al. (2025) automates the discovery of optimal configurations for agents which successive halving to explore the space of agentic and non-agentic prompting patterns. The sequential optimal learning approach for automated prompt engineering Wang et al. (2025) uses Bayesian regression and Knowledge-Gradient policies to efficiently identify effective prompt features. Progressively Automatic Prompt Optimization Qu et al. (2025) introduces an evolution-based algorithm to optimize prompts for visual classification tasks.
@@
We propose a hybrid framework integrating LLM-driven rewriting with natural language feedback Pryzant et al. (2023), alongside self-reflection Shinn et al. (2024) and planning Wang et al. (2023), enhancing prompt adaptability and precision.
@@
Data Synthesis: Using large language models (LLMs) for data synthesis is a relatively new and rapidly evolving approach. Recent advancements have shown that LLMs possess the capability to generate text with fluency and quality comparable to human output Li et al. (2023); Mukherjee et al. (2023); Eldan and Li (2023). For instance, prior work Gao et al. (2022) has explored leveraging pre-trained language models (PLMs) to generate task-specific text data that can be used to train and evaluate. Recent work Magpie Xu et al. (2024) leverages the auto-regressive nature of aligned LLMs to generate high-quality instruction data. Additionally, Synthetic Text Generation for Training Large Language Models via Gradient Matching Nguyen et al. (2025) proposes a novel approach to generate synthetic text that matches the gradients of human data. However, these studies have not fully incorporated advanced methodologies such as chain-of-thought (CoT) reasoning, in-context learning, or data synthesis driven by prompts that integrate task descriptions and label information.
@@
In this study, we systematically experimented with a range of techniques, including in-context learning and prompt-driven data synthesis, combining task descriptions and label information. Our results indicate that integrating these approaches produces high-quality synthetic data. To further enhance the robustness and applicability of the synthetic data, we introduced a difficulty tier, making the generated data more challenging. These findings highlight the potential of combining advanced LLM capabilities with tailored prompting strategies to improve data synthesis quality and reliability for prompt optimization.
@@ 3 Method
SIPDO presents a two-agent system for optimizing prompts using data augmentation techniques. The workflow has two cooperating agents: (i) Data Generator creates synthetic data with increasing difficulty levels to expose weaknesses in the prompt, and (ii) Auto Prompt Optimizer iteratively analyzes errors and rewrites the prompt to maximize task performance. The two agents advance in lock-step, and together they grow a prompt that remains compact across increasingly challenging samples .An overview of SIPDO is shown in Fig 1.
Notation.
We define the true data distribution as $S$, which governs input-label pairs $(x,y)\in\mathcal{X}\times\mathcal{Y}$. Let $N$ denote the size of an i.i.d. dataset drawn from $S$, denoted as $\{(x_{i},y_{i})\}_{i=1}^{N}\sim S$. We consider a fixed large language model equipped with a prompt $p\in\mathcal{P}$, and define its output function as $f(p,x)\in\mathcal{Y}$. Prediction accuracy is measured using a bounded surrogate loss $L\bigl{(}f(p,x),y\bigr{)}$, where $L\in[0,1]$.
To help improve the quality of prompts, we introduce a synthetic data generator defined by a distribution $q_{\psi}(\tilde{x},\tilde{y})$, parameterized by $\psi\in\Psi$, which produces synthetic samples forming a dataset $D=\{(\tilde{x}_{i},\tilde{y}_{i})\}_{i=1}^{M}$, where $M$ is the number of generated examples. To ensure that the synthetic labels remain realistic, we estimate the population label prior with $p^{\ast}(y)$, and use this to regularize the generator.
@@ 3.1 Data Generator
Sampling rule.
The data generator first draws a target label $\tilde{y}\sim p^{*}(y)$. By sampling a latent variable $z\sim g_{\phi}(z|S)$ that captures the structure of few-shot $S$, the decoder $q_{\psi}$ produces $\tilde{x}=q_{\psi}(z,\tilde{y},c)$ where c is a controlled difficulty tier.
Learning objective.
The parameters $\psi$ are learned by minimizing a hybrid objective that balances the KL penalty and the bounded surrogate loss:
@@
Progressive complexity.
To address tasks of varying complexity, we introduces a progressive complexity parameter $c$ where $c\in\{1,...,n\}$ so that prompts could be tested on gradually more challenging examples. This allows the prompts to progressively improve and generalize effectively across task of increasing difficulty. Since $q_{\psi}$ is conditioned on $c$, a single latent template $(z,y)$ can therefore yield $n$ difficulty-aligned variants
@@
For curriculum generation, an ordered sequence $c_{1}<\dots<c_{L}$ is sampled and feeds the output of the previous level back into the generator,
$\tilde{x}^{(1)}=q_{\psi}\bigl{(}z,\,y,\,c_{1}\bigr{)},$
$\tilde{x}^{(2)}=q_{\psi}\bigl{(}h_{\phi}\bigl{(}x^{(1)}\bigr{)},\,y,\,c_{2}\bigr{)},$
$\vdots$
$\tilde{x}^{(L)}=q_{\psi}\bigl{(}h_{\phi}\bigl{(}x^{(L-1)}\bigr{)},\,y,\,c_{L}\bigr{)}.$
where $h_{\phi}$ is a summarizer that distills the previous sample into a new latent cue, allowing semantic depth to accumulate across levels. The sequence $c_{1}<c_{2}<\cdots<c_{L}$ guarantees monotone growth of problem complexity, providing a rich gradient of difficulty for the prompt to learn from.
@@ 3.2 Auto Prompt Optimizer
After each new synthetic instance is calibrated, the Auto Prompt Optimizer probes the current prompt, identifies the weaknesses, and repairs them before the next instance is drawn. This stage avoids hard-to-diagnose failures and builds a prompt that is both robust and suitable for specific tasks.
Accuracy score.
At iteration $t\in\{1,\dots,M\}$, the optimizer improves the current prompt $p^{(t)}$ using the feedback collected from synthetic log $D_{t}=\{(\tilde{x}_{j},\tilde{y}_{j})\}^{t}_{j=1}\subseteq\mathcal{X}\times\mathcal{Y}$. For any prompt $p$ and set $\mathcal{A}\subseteq\mathcal{X}\times\mathcal{Y}$, we define
@@
Step 1: Error analysis.
We first evaluate $p^{(t)}$ on the whole set and collect the current error slice
@@
If $\mathcal{E}^{(t)}=\varnothing$, the prompt already "covers" all unseen cases, therefore, we terminate and return $p^{*}=p^{(t)}$; otherwise, we proceed to the next step.
Step 2: Recommendation.
A reflection module $\mathcal{R}_{\varphi}$ inspects $\mathcal{E}^{(t)}$ and produces a textual-patch suggesting how to modify the prompt:
$\Delta^{(t)}=\mathcal{R}_{\varphi}\!\bigl{(}p^{(t)},\mathcal{E}^{(t)}\bigr{)}.$
This summarizes why the prompt failed and how it can be amended(e.g., add boundary cases, clarify/revise instructions, drop distracting details).
Step 3: Targeted refinement.
A prompt editor $\mathcal{U}_{\theta}$ applies the patch $\Delta^{(t)}$ to a revised prompt $\tilde{p}^{(t)}$ in order to fix the current error
$\tilde{p}^{(t)}=\mathcal{U}_{\theta}\!\bigl{(}\Delta^{(t)},p^{(t)},\mathcal{E}^{(t)}\bigr{)}.$
Local confirmation.
We then test revised prompt $\tilde{p}^{(t)}$ only on the current errors: if $s_{\mathcal{E}^{(t)}}(\tilde{p}^{(t)})<1$, some errors still remain. In this case, we make the revised prompt as new baseline prompt by setting $p^{(t)}\leftarrow\tilde{p}^{(t)}$, updating $\mathcal{E}^{(t)}$, and repeating Step 2 to generate more sufficient patch $\Delta^{(t)}$; otherwise, proceed to global confirmation.
Global confirmation.
Solving the local error slice is not enough-we must ensure that revised prompt "covers" all seen cases. Therefore, we evaluate $\tilde{p}^{(t)}$ on the entire synthetic history collected seen so far by $s_{{D}_{t}}\!\bigl{(}\tilde{p}^{(t)}\bigr{)}$. During evaluation, if $\mathcal{E}^{(t)}\neq\varnothing$ at any previous data, we treat them as new error set and sent them back to step 2 with new $\mathcal{E}^{(t)}$ to fix the current error. If $\mathcal{E}^{(t)}=\varnothing$, we accept the revision, set $p^{(t+1)}=\tilde{p}^{(t)}$, draw the next synthetic example, and restart from Step 1 until $t=M$.
Convergence guarantee.
Because $s_{D}(p^{(t)})$ is non-decreasing and bounded above by 1, the process stops at most $M$ successful corrections or the user-chosen cap $T_{\max}$. The final output
$p^{*}\;=\;\arg\max_{0\leq t\leq T}s_{{D}}\!\bigl{(}p^{(t)}\bigr{)}$
achieves perfect coverage $(s_{{D}_{T}}(p^{*})=1)$ whenever it is attainable within the budget.
This revised loop mirrors practical prompt-debugging: it first addresses specific failure case, then confirms that the updated prompt continues to perform correctly on all previously solved examples. By iteratively applying this feedback-driven process, the system systematically refines prompts to improve clarity, adaptability, and overall performance, making the framework highly generalizable across diverse tasks and domains.
@@ 3.3 Theoretical Guarantee
Since one of our goals in SIPDO is to demonstrate that, data augmentation, a popular branch of performance improvement in deep learning, can also be used in prompt optimization context, we aim to offer similiar performance guarantees as done in previous data augmentation literature Wang et al. (2022); Chen et al. (2020); Dao et al. (2019).
Assumptions.
We first offer the assumptions that we need for the theoretical guarantees.
A1 (Label-preservation)
@@
We require the generator never flips the ground-truth label of the base example it is derived from (it may, however, hallucinate novel inputs as long as their labels match the intended classes).
@@
Let
$\psi^{\star}=\arg\max_{\psi\in\Psi}\mathbb{E}_{q_{\psi}}L(f(p,\tilde{x}),\tilde{y})-\lambda^{-1}R(\psi),$
The inner-loop training of the generator attains a value at most $\varepsilon$ below this supremum.
@@
Wang et al. (2022) For every prompt $p$, the empirical loss deviates from its population counterpart by at most $q(|\mathcal{P}|,n,\delta)$ with probability $1-\delta$., where a standard form of $q(|\mathcal{P}|,n,\delta)$ is $\tilde{O}\left(\sqrt{\frac{\log|\mathcal{P}|+\log(1/\delta)}{N}}\right).$
@@
The KL penalty controls how far the generator may wander: if it manufactures rare-label outliers, $R(\psi)$ increases and the bound tightens. We can verify that $q_{\psi}(y)$ is always absolutely-continuous w.r.t. $p^{*}(y)$; KL is then finite and the inequality follows from the classical Donsker–Varadhan variational formula.
@@
The 001111 loss is upper-bounded by the surrogate loss:
$\mathbbm{1}\{f(p,x)\neq y\}\leq L(f(p,x),y).$
@@
Theorem 3.1 Regularised Worst-case Data Generation
Under Assumptions B1, B2, B4–B6, for any fixed prompt $p\in\mathcal{P}$, with probability at least $1-\delta$ over the draw of the training set, we have
$\begin{split}\underbrace{\sup_{\psi\in\Psi}\mathbb{E}_{q_{\psi}}\mathbbm{1}\{f(p,\tilde{x})\neq\tilde{y}\}}_{\begin{subarray}{c}\text{population}\\ \text{worst-case error}\end{subarray}}&\;\leq\;\underbrace{\tfrac{1}{n}\sum_{i=1}^{n}L\bigl{(}f(p,x_{i}),y_{i}\bigr{)}}_{\text{empirical risk}}\\ &\quad+\underbrace{\lambda^{-1}R(\psi^{\star})}_{\begin{subarray}{c}\text{KL penalty of}\\ \text{hardest generator}\end{subarray}}\\ &\quad+\;\varepsilon+\;q\bigl{(}|\mathcal{P}|,n,\delta\bigr{)}.\end{split}$ (2)
Practical implication
The inequality states that if the empirical loss of the prompt is low, and no generator can inflate that loss without paying a high KL tax, then even a hypothetically all-powerful adversary (generator) cannot cause the prompt to misclassify more than the RHS. Selecting a larger $\lambda$ tightens the KL tax, thus lowering the worst-case error but potentially harming accuracy—precisely the robustness–performance trade-off observed empirically in Section 4.

### v3

@@ 1 Introduction
Large language models (LLMs) have demonstrated strong performance across a wide range of natural language tasks, including classification, question answering, and reasoning. However, their output quality is highly sensitive to prompt design—small changes in phrasing, structure, or formatting can lead to significant variations in performance (He et al., 2024; Spiess et al., 2025). This sensitivity has made prompt optimization a core challenge in adapting LLMs to downstream applications, where consistency and reliability are crucial. In domains such as healthcare and finance, the inability to ensure stable, predictable performance makes prompt optimization more than a desirable enhancement, but a critical necessity for reliable system deployment.
The core challenge of prompt optimization is multifaceted. Unlike traditional hyperparameter tuning, the search space of prompts is discrete, non-differentiable, and vast. Small modifications to prompts can have unpredictable effects on LLMs’ behavior, and the gradient-based methods that typically power optimization in machine learning are not directly applicable. Furthermore, LLMs often perform well on a fixed, curated test set, but their performance can deteriorate when faced with novel linguistic variations, edge cases, or adversarial queries. This makes prompt optimization particularly challenging, as a prompt that performs well in one scenario may fail when the input distribution shifts, leading to issues such as catastrophic forgetting or fragile performance across different contexts.
@@
Prior work in prompt optimization has explored manual tuning, discrete search, and gradient-based methods to improve model responses (Wang et al., 2023; Shin et al., 2020; Cui et al., 2024; Kwon et al., 2024; Zhang et al., 2024). While effective in some settings, these approaches do not address the dynamic nature of real-world inputs, where the input space evolves over time. As a result, they can produce prompts that perform well on average but lack robustness when the input distribution changes.
In contrast, data augmentation has long been used in supervised learning to improve model robustness by exposing models to diverse training conditions (Mikołajczyk & Grochowski, 2018). In the context of prompt learning, the ability of LLMs to generate high-quality synthetic data presents an exciting opportunity to improve prompt optimization. However, existing prompt optimization methods do not leverage synthetic data in a dynamic, feedback-driven manner (Singh et al., 2023; Gilardi et al., 2023; Tang et al., 2023; Gao et al., 2023). Moreover, it is not sufficient to simply produce more data; the challenge is to generate data that is purposeful and stressful that targets the current failure modes of the prompt and provides a progressive challenge that helps guide its evolution. The synthetic data must also be carefully crafted to ensure that it does not overwhelm the model with trivially easy or overly difficult cases, but instead exposes latent weaknesses that need to be addressed.
@@
To address these challenges, we propose SIPDO (Self-Improving Prompts through Data-Augmented Optimization), a closed-loop framework for prompt optimization that integrates synthetic data generation directly into the learning process. SIPDO consists of two components: a Synthetic Data Generator that produces inputs specifically designed to challenge the current prompt, and a Prompt Optimizer that uses these examples to iteratively refine the prompt. This feedback loop enables the prompt to evolve continuously over time, adapting to new, previously unseen inputs without requiring external supervision or the need to tune for each new scenario individually. Particularly, SIPDO addresses the unique challenges of prompt optimization by transforming the optimization process from a static, one-time procedure to a dynamic, self-adaptive learning loop. This shift is critical for ensuring prompt robustness in the face of evolving input distributions.
@@
  • We empirically demonstrate that augmenting prompt optimization with synthetic data improves performance across multiple reasoning benchmarks, surpassing existing prompt tuning methods.
@@ 2.1 Automatic Prompt Engineering
Automatically discovering optimal prompts has become a key challenge in the era of large language models (LLMs). Automatic Prompt Engineering (APE) employs optimization-based, generative, and template-driven approaches. Optimization techniques include gradient-based search (Shin et al., 2020), reinforcement learning (Ouyang et al., 2022; Kwon et al., 2024), and evolutionary algorithms (Cui et al., 2024). Generative methods use models like GPT and Gemini to generate candidate prompts, with StablePrompt (Kwon et al., 2024) optimizing prompts via reinforcement learning. Additionally, PromptAgent (Wang et al., 2023) breaks down prompt creation into sub-goals, while template-driven approaches, like fill-in-the-blank formats, ensure clarity (Chen et al., 2024). Recent work has expanded on automatic prompt optimization techniques. AutoPDL (Spiess et al., 2025) automates the discovery of optimal configurations for agents which successive halving to explore the space of agentic and non-agentic prompting patterns. The sequential optimal learning approach for automated prompt engineering (Wang et al., 2025) uses Bayesian regression and Knowledge-Gradient policies to efficiently identify effective prompt features. Progressively Automatic Prompt Optimization (Qu et al., 2025) introduces an evolution-based algorithm to optimize prompts for visual classification tasks.
Adopting LLM as a feedback loop to refine prompts has recently emerged. Self-Refine (Madaan et al., 2023) improves the model by generating feedback based on the previously generated outputs and using that feedback to refine the next output. In Promptbreeder (Fernando et al., 2023), it uses an iterative feedback loop to select better prompts from the original prompts and mutation prompts, while Recursive In-Context Learning for Autonomous Prompt Generation in Large Language Models: A Self-Instructed Approach (Yilar et al., 2024) introduces a framework that refines prompts through iterative loops based on generated outputs. More recently, SPO(Self-Supervised Prompt Optimization) (Xiang et al., 2025) uses iterative feedback to refine prompts by comparing the outputs of the current prompt and its revised version without relying on ground truth, and DLPO (Peng et al., 2025) optimizes prompts within loops by mimicking a deep learning style where it uses textual loss, gradients, and a feedback process to update prompts. CriSPO (He et al., 2025) also extends iterative feedback loop optimization using critique guides to update prompts with multi-metric.
@@
We propose a hybrid framework integrating LLM-driven rewriting with natural language feedback (Pryzant et al., 2023), alongside self-reflection (Shinn et al., 2024) and planning (Wang et al., 2023), enhancing prompt adaptability and precision.
2.2 Data Synthesis
@@
Using large language models (LLMs) for data synthesis is a relatively new and rapidly evolving approach. Recent advancements have shown that LLMs possess the capability to generate text with fluency and quality comparable to human output (Li et al., 2023; Mukherjee et al., 2023; Eldan & Li, 2023). For instance, prior work (Gao et al., 2023) has explored leveraging pre-trained language models (PLMs) to generate task-specific text data that can be used to train and evaluate. Recent work Magpie (Xu et al., 2024) leverages the auto-regressive nature of aligned LLMs to generate high-quality instruction data. Additionally, Synthetic Text Generation for Training Large Language Models via Gradient Matching (Nguyen et al., 2025) proposes a novel approach to generate synthetic text that matches the gradients of human data. However, these studies have not fully incorporated advanced methodologies such as chain-of-thought (CoT) reasoning, in-context learning, or data synthesis driven by prompts that integrate task descriptions and label information.
@@
In this study, we systematically experimented with a range of techniques, including in-context learning and prompt-driven data synthesis, combining task descriptions and label information. Our results show that these approaches generate high-quality synthetic data. By introducing a difficulty tier, we further enhanced the data’s robustness and applicability. These findings demonstrate the potential of combining advanced LLM capabilities with tailored prompting strategies to improve data synthesis for prompt optimization.
@@ 3 Method
In this work, we introduce SIPDO, a two-agent system for optimizing prompts using data augmentation techniques. The workflow has two cooperating agents: (i) Data Generator creates synthetic data with increasing difficulty levels to expose weaknesses in the prompt, and (ii) Auto Prompt Optimizer iteratively analyzes errors and rewrites the prompt to maximize task performance. An overview of SIPDO is shown in Fig 1.
[…]
@@ 3.1 Data Generator
Sampling rule. The data generator first draws a target label $\tilde{y}\sim p^{*}(y)$. By sampling a latent variable $z\sim g_{\phi}(z|S)$ that captures the structure of few-shot $S$, the decoder $q_{\psi}$ produces $\tilde{x}=q_{\psi}(z,\tilde{y},c)$ where c is a controlled difficulty tier.
Learning objective. The parameters $\psi$ are learned by minimizing a hybrid objective that balances the KL penalty and the bounded surrogate loss:
@@
Progressive difficulty. To address tasks of varying difficulty, we introduces a progressive difficulty parameter $c$ where $c\in\{1,...,n\}$ so that prompts could be tested on gradually more challenging examples. This allows the prompts to progressively improve and generalize effectively across task of increasing difficulty. Since $q_{\psi}$ is conditioned on $c$, a single latent template $(z,y)$ can therefore yield $n$ difficulty-aligned variants
@@
For curriculum generation, an ordered sequence $c_{1}<\dots<c_{n}$ is sampled and feeds the output of the previous level back into the generator,
$\tilde{x}^{(1)}=q_{\psi}\bigl(z,\,y,\,c_{1}\bigr),\tilde{x}^{(2)}=q_{\psi}\bigl(h_{\phi}\bigl(x^{(1)}\bigr),\,y,\,c_{2}\bigr),\dots,\tilde{x}^{(n)}=q_{\psi}\bigl(h_{\phi}\bigl(x^{(n-1)}\bigr),\,y,\,c_{n}\bigr),$
where $h_{\phi}$ is a summarizer that distills the previous sample into a new latent cue, allowing semantic depth to accumulate across levels. The sequence $c_{1}<c_{2}<\cdots<c_{n}$ guarantees monotone growth of problem difficulty, providing a rich gradient of difficulty for the prompt to learn from.
@@ 3.2 Auto Prompt Optimizer
After each new synthetic instance is calibrated, the Auto Prompt Optimizer probes the current prompt, identifies the weaknesses, and repairs them before the next instance is drawn. This stage builds a prompt that is both robust, suitable, and generalizable for specific tasks.
Accuracy score. At iteration $t\in\{1,\dots,M\}$, the optimizer improves the current prompt $p^{(t)}$ using the feedback collected from synthetic log $D_{t}=\{(\tilde{x}_{j},\tilde{y}_{j})\}^{t}_{j=1}\subseteq\mathcal{X}\times\mathcal{Y}$. For any prompt $p$ and set $\mathcal{A}\subseteq\mathcal{X}\times\mathcal{Y}$, we define
@@
Step 1: Error analysis. We first evaluate $p^{(t)}$ on the whole set and collect the current error slice
@@
If $\mathcal{E}^{(t)}=\varnothing$, the prompt already “covers” all unseen cases, therefore, we terminate and return $p^{*}=p^{(t)}$; otherwise, we proceed to the next step.
[…]
Step 3: Targeted refinement. A prompt editor $\mathcal{U}_{\theta}$ applies the patch $\Delta^{(t)}$ to a revised prompt $\tilde{p}^{(t)}$ in order to fix the current error: $\tilde{p}^{(t)}=\mathcal{U}_{\theta}\!\bigl(\Delta^{(t)},p^{(t)},\mathcal{E}^{(t)}\bigr).$
Local confirmation. We then test revised prompt $\tilde{p}^{(t)}$ only on the current errors: if $s_{\mathcal{E}^{(t)}}(\tilde{p}^{(t)})<1$, some errors still remain. In this case, we make the revised prompt as new baseline prompt by setting $p^{(t)}\leftarrow\tilde{p}^{(t)}$, updating $\mathcal{E}^{(t)}$, and repeating Step 2 to generate more sufficient patch $\Delta^{(t)}$; otherwise, proceed to global confirmation.
Global confirmation. Solving the local error slice is not enough-we must ensure that revised prompt “covers” all seen cases. Therefore, we evaluate $\tilde{p}^{(t)}$ on the entire synthetic history collected seen so far by $s_{{D}_{t}}\!\bigl(\tilde{p}^{(t)}\bigr)$. During evaluation, if $\mathcal{E}^{(t)}\neq\varnothing$ at any previous data, we treat them as new error set and sent them back to step 2 with new $\mathcal{E}^{(t)}$ to fix the current error. If $\mathcal{E}^{(t)}=\varnothing$, we accept the revision, set $p^{(t+1)}=\tilde{p}^{(t)}$, draw the next synthetic example, and restart from Step 1 until $t=M$.
Convergence guarantee. Because $s_{D}(p^{(t)})$ is non-decreasing and bounded above by 1, the process stops at most $M$ successful corrections or the user-chosen cap $T_{\max}$. The final output $p^{*}\;=\;\arg\max_{0\leq t\leq T}s_{{D}}\!\bigl(p^{(t)}\bigr)$ achieves perfect coverage $(s_{{D}_{T}}(p^{*})=1)$ whenever it is attainable within the budget.
By iteratively applying this feedback-driven process, it systematically refines prompts to improve clarity, adaptability, and overall performance, making the framework highly generalizable across tasks and domains.
@@ 3.3 Theoretical Guarantee
Since one of our goals in SIPDO is to demonstrate that, data augmentation, a popular branch of performance improvement in deep learning, can also be used in prompt optimization context, we aim to offer similiar performance guarantees as done in previous data augmentation literature (Wang et al., 2022; Chen et al., 2020; Dao et al., 2019).
Assumptions. We first offer the assumptions that we need for the theoretical guarantees.
A1 (Label‑preservation)
@@
We require the generator never flips the ground‑truth label of the base example it is derived from (it may, however, hallucinate novel inputs as long as their labels match the intended classes). In practice, because LLMs sometimes assign unexpected labels, we first generate the label $\tilde{y}$ and then sample $\tilde{x}$ conditioned on that label. For tasks and domains where producing valid synthetic data is difficult, we apply a three-voter check: three expert agents independently verify each generated item for label–input consistency and basic factual correctness.
@@
Let $\psi^{\star}=\arg\max_{\psi\in\Psi}\mathbb{E}_{q_{\psi}}L(f(p,\tilde{x}),\tilde{y})-\lambda^{-1}R(\psi),$ The inner‑loop training of the generator attains a value at most $\varepsilon$ below this supremum.
@@
(Wang et al., 2022) For every prompt $p$, the empirical loss deviates from its population counterpart by at most $q(|\mathcal{P}|,n,\delta)$ with probability $1-\delta$., where a standard form of $q(|\mathcal{P}|,n,\delta)$ is $\tilde{O}\left(\sqrt{\frac{\log|\mathcal{P}|+\log(1/\delta)}{N}}\right).$
@@
The KL penalty controls how far the generator may wander: if it manufactures rare‑label outliers, $R(\psi)$ increases and the bound tightens. We can verify that $q_{\psi}(y)$ is always absolutely-continuous w.r.t. $p^{*}(y)$; KL is then finite and the inequality follows from the classical Donsker–Varadhan variational formula.
@@
The 011 loss is upper‑bounded by the surrogate loss: $\mathbbm{1}\{f(p,x)\neq y\}\leq L(f(p,x),y).$
@@
Theorem 3.1 Regularised Worst-case Data Generation. Under Assumptions A1-A5, for any fixed prompt $p\in\mathcal{P}$, with probability at least $1-\delta$ over the draw of the training set, we have
$\underbrace{\sup_{\psi\in\Psi}\mathbb{E}_{q_{\psi}}\mathbbm{1}\{f(p,\tilde{x})\neq\tilde{y}\}}_{\begin{subarray}{c}\text{population}\\ \text{worst-case error}\end{subarray}}\;\leq\;\underbrace{\tfrac{1}{n}\sum_{i=1}^{n}L\bigl(f(p,x_{i}),y_{i}\bigr)}_{\text{empirical risk}}\\ \quad+\underbrace{\lambda^{-1}R(\psi^{\star})}_{\begin{subarray}{c}\text{KL penalty of}\\ \text{hardest generator}\end{subarray}}\\ \quad+\;\varepsilon+\;q\bigl(|\mathcal{P}|,n,\delta\bigr).$ (1)
Practical implication. The inequality states that if the empirical loss of the prompt is low, and no generator can inflate that loss without paying a high KL tax, then even a hypothetically all‑powerful adversary (generator) cannot cause the prompt to misclassify more than the RHS. Selecting a larger $\lambda$ tightens the KL tax, thus lowering the worst‑case error but potentially harming accuracy—precisely the robustness–performance trade‑off observed empirically in Experiments Section 4. In addition, detailed proof of theorem can be found in Appendix F.
MMLU Subject Method 1st iteration 2nd iteration 3rd iteration Final Result (Comparative Acc.)
College Computer Science TextGrad 85 88 86 89($\downarrow 4.0$)
M-TextGrad 85 87 87 89($\downarrow 4.0$)
REVOLVE 85 88 89 90($\downarrow 3.0$)
SIPDO 93
Machine Learning TextGrad 84.8 87.5 82.1 88.4($\downarrow 5.4$)
M-TextGrad 85.5 85.4 85.3 85.0($\downarrow 8.8$)
REVOLVE 85.7 86.6 85.7 88.4($\downarrow 5.4$)
ANN 90.1($\downarrow 3.7$)
SIPDO 93.8
College Biology TextGrad 95.1 97.2 95.1 96.5($\downarrow 0.0$)
REVOLVE 96.5 96.1 97.2 96.5($\downarrow 0.0$)
SIPDO 96.5
Table 1: Results on MMLU Machine Learning, College Computer Science, and College Biology subject by GPT-4o, demonstrating SIPDO’s effectiveness on different subjects